# Provably Mitigating Overoptimization in RLHF: Your SFT Loss is Implicitly an Adversarial Regularizer

### v1

@@ A.2 Technical Lemmas
$\mathcal{L}_{\mathcal{D}}(r^{\star})-\mathcal{L}_{\mathcal{D}}(r)$ $\leq-2\cdot\mathbb{E}_{(x,a^{1},a^{0})\sim\mu_{\mathcal{D}}(\cdot,\cdot,\cdot)}\Big{[}D_{\mathrm{Hellinger}}^{2}\big{(}\mathbb{P}_{r^{\star}}(\cdot|x,a^{1},a^{0})\|\mathbb{P}_{r}(\cdot|x,a^{1},a^{0})\big{)}\Big{]}$ (A.47)
$\qquad+\frac{3}{N}\cdot\log\left(\frac{\mathcal{N}_{\varepsilon}(\mathcal{R},\|\cdot\|_{\infty})}{\delta}\right).$ (A.48)
Proof of Lemma A.1.
@@
$\mathcal{L}_{\mathcal{D}}(r^{\star})-\mathcal{L}_{\mathcal{D}}(r)$ $\leq-2\cdot\mathbb{E}_{(x,a^{1},a^{0})\sim\mu_{\mathcal{D}}(\cdot,\cdot,\cdot)}\Big{[}D_{\mathrm{Hellinger}}^{2}\big{(}\mathbb{P}_{r^{\star}}(\cdot|x,a^{1},a^{0})\|\mathbb{P}_{r}(\cdot|x,a^{1},a^{0})\big{)}\Big{]}$ (A.49)
$\qquad+\frac{2}{N}\cdot\log\left(\frac{\mathcal{N}_{\varepsilon}(\mathcal{R},\|\cdot\|_{\infty})}{\delta}\right).$ (A.50)
@@
$\mathcal{L}_{\mathcal{D}}(r^{\star})-\mathcal{L}_{\mathcal{D}}(r)$ $=\mathcal{L}_{\mathcal{D}}(r^{\star})-\mathcal{L}_{\mathcal{D}}(r^{\dagger})+\mathcal{L}_{\mathcal{D}}(r^{\dagger})-\mathcal{L}_{\mathcal{D}}(r)$ (A.51)
$\leq-2\cdot\mathbb{E}_{(x,a^{1},a^{0})\sim\mu_{\mathcal{D}}(\cdot,\cdot,\cdot)}\Big{[}D_{\mathrm{Hellinger}}^{2}\big{(}\mathbb{P}_{r^{\star}}(\cdot|x,a^{1},a^{0})\|\mathbb{P}_{r^{\dagger}}(\cdot|x,a^{1},a^{0})\big{)}\Big{]}$ (A.52)
$\qquad+\frac{2}{N}\cdot\log\left(\frac{\mathcal{N}_{\varepsilon}(\mathcal{R},\|\cdot\|_{\infty})}{\delta}\right)+\mathcal{L}_{\mathcal{D}}(r^{\dagger})-\mathcal{L}_{\mathcal{D}}(r)$ (A.53)
$\leq-2\cdot\mathbb{E}_{(x,a^{1},a^{0})\sim\mu_{\mathcal{D}}(\cdot,\cdot,\cdot)}\Big{[}D_{\mathrm{Hellinger}}^{2}\big{(}\mathbb{P}_{r^{\star}}(\cdot|x,a^{1},a^{0})\|\mathbb{P}_{r}(\cdot|x,a^{1},a^{0})\big{)}\Big{]}$ (A.54)
$\qquad+\frac{2}{N}\cdot\log\left(\frac{\mathcal{N}_{\varepsilon}(\mathcal{R},\|\cdot\|_{\infty})}{\delta}\right)+\mathcal{L}_{\mathcal{D}}(r^{\dagger})-\mathcal{L}_{\mathcal{D}}(r)$ (A.55)
$\qquad\qquad+4\cdot\mathbb{E}_{(x,a^{1},a^{0})\sim\mu_{\mathcal{D}}(\cdot,\cdot,\cdot)}\Big{[}D_{\mathrm{Hellinger}}^{2}\big{(}\mathbb{P}_{r^{\dagger}}(\cdot|x,a^{1},a^{0})\|\mathbb{P}_{r}(\cdot|x,a^{1},a^{0})\big{)}\Big{]},$ (A.56)
where in the fir inequality we use (A.50) for $r^{\dagger}$ and in the second inequality we utilize the triangular inequality for Hellinger distance. Therefore, it remains to upper bound the approximation error induced by $r^{\dagger}$. On the one hand, by the definition of $\mathcal{L}_{\mathcal{D}}$ in (3.2), we have that
$\mathcal{L}_{\mathcal{D}}(r^{\dagger})-\mathcal{L}_{\mathcal{D}}(r)$ $=\frac{1}{N}\sum_{i=1}^{N}y_{i}\cdot\log\left(\frac{\sigma\big{(}r(x_{i},a^{1}_{i})-r(x_{i},a^{0}_{i})\big{)}}{\sigma\big{(}r^{\dagger}(x_{i},a^{1}_{i})-r^{\dagger}(x_{i},a^{0}_{i})\big{)}}\right)$ (A.57)
$\qquad+\frac{1}{N}\sum_{i=1}^{N}(1-y_{i})\cdot\log\left(\frac{\sigma\big{(}r(x_{i},a^{0}_{i})-r(x_{i},a^{1}_{i})\big{)}}{\sigma\big{(}r^{\dagger}(x_{i},a^{0}_{i})-r^{\dagger}(x_{i},a^{1}_{i})\big{)}}\right).$ (A.58)
@@
$\mathcal{L}_{\mathcal{D}}(r^{\dagger})-\mathcal{L}_{\mathcal{D}}(r)$ $\leq\frac{1}{N}\sum_{i=1}^{N}y_{i}\cdot\frac{\sigma\big{(}r(x_{i},a^{1}_{i})-r(x_{i},a^{0}_{i})\big{)}-\sigma\big{(}r^{\dagger}(x_{i},a^{1}_{i})-r^{\dagger}(x_{i},a^{0}_{i})\big{)}}{\sigma\big{(}r^{\dagger}(x_{i},a^{1}_{i})-r^{\dagger}(x_{i},a^{0}_{i})\big{)}}$ (A.59)
$\qquad\qquad+\frac{1}{N}\sum_{i=1}^{N}(1-y_{i})\cdot\frac{\sigma\big{(}r(x_{i},a^{0}_{i})-r(x_{i},a^{1}_{i})\big{)}-\sigma\big{(}r^{\dagger}(x_{i},a^{0}_{i})-r^{\dagger}(x_{i},a^{1}_{i})\big{)}}{\sigma\big{(}r^{\dagger}(x_{i},a^{0}_{i})-r^{\dagger}(x_{i},a^{1}_{i})\big{)}}.$ (A.60)
Now since $\|r^{\dagger}-r\|_{\infty}\leq\varepsilon$ and $r^{\dagger}\in[0,R]$, invoking Lemma A.2, we can derive that
$\mathcal{L}_{\mathcal{D}}(r^{\dagger})-\mathcal{L}_{\mathcal{D}}(r)$ $\leq\frac{1}{N}\sum_{i=1}^{N}\frac{\big{|}\big{(}r(x_{i},a^{1}_{i})-r(x_{i},a^{0}_{i})\big{)}-\big{(}r^{\dagger}(x_{i},a^{1}_{i})-r^{\dagger}(x_{i},a^{0}_{i})\big{)}\big{|}}{(1+e^{R})^{-1}}$ (A.61)
$\qquad+\frac{1}{N}\sum_{i=1}^{N}\frac{\big{|}\big{(}r(x_{i},a^{0}_{i})-r(x_{i},a^{1}_{i})\big{)}-\big{(}r^{\dagger}(x_{i},a^{0}_{i})-r^{\dagger}(x_{i},a^{1}_{i})\big{)}\big{|}}{(1+e^{R})^{-1}}$ (A.62)
$\leq 4\cdot\|r^{\dagger}-r\|_{\infty}\cdot(1+e^{R})\leq 4\varepsilon\cdot(1+e^{R}).$ (A.63)
@@
$D_{\mathrm{Hellinger}}^{2}\big{(}\mathbb{P}_{r^{\dagger}}(\cdot|x,a^{1},a^{0})\|\mathbb{P}_{r}(\cdot|x,a^{1},a^{0})\big{)}$ (A.64)
$\qquad\leq D_{\mathrm{TV}}\big{(}\mathbb{P}_{r^{\dagger}}(\cdot|x,a^{1},a^{0})\|\mathbb{P}_{r}(\cdot|x,a^{1},a^{0})\big{)}$ (A.65)
$\qquad=\Big{|}\sigma\big{(}r^{\dagger}(x,a^{1})-r^{\dagger}(x,a^{0})\big{)}-\sigma\big{(}r(x,a^{1})-r(x,a^{0})\big{)}\Big{|}$ (A.66)
$\qquad\leq\Big{|}\big{(}r^{\dagger}(x,a^{1})-r^{\dagger}(x,a^{0})\big{)}-\big{(}r(x,a^{1})-r(x,a^{0})\big{)}\Big{|}$ (A.67)
$\qquad\leq 2\cdot\|r^{\dagger}-r\|_{\infty}\leq 2\varepsilon,$ (A.68)
where the first inequality uses the fact that $D_{\mathrm{Hellinger}}^{2}\leq D_{\mathrm{TV}}$, the equality uses the same argument as (A.35), and the second inequality applies Lemma A.2. Finally, combining (A.56), (A.63), and (A.68), we conclude that
$\mathcal{L}_{\mathcal{D}}(r^{\star})-\mathcal{L}_{\mathcal{D}}(r)$ $\leq-2\cdot\mathbb{E}_{(x,a^{1},a^{0})\sim\mu_{\mathcal{D}}(\cdot,\cdot,\cdot)}\Big{[}D_{\mathrm{Hellinger}}^{2}\big{(}\mathbb{P}_{r^{\star}}(\cdot|x,a^{1},a^{0})\|\mathbb{P}_{r}(\cdot|x,a^{1},a^{0})\big{)}\Big{]}$ (A.69)
$\qquad+\frac{2}{N}\cdot\log\left(\frac{\mathcal{N}_{\varepsilon}(\mathcal{R},\|\cdot\|_{\infty})}{\delta}\right)+6\varepsilon\cdot(1+e^{R}).$ (A.70)
@@
$\mathcal{L}_{\mathcal{D}}(r^{\star})-\mathcal{L}_{\mathcal{D}}(r)$ $\leq-2\cdot\mathbb{E}_{(x,a^{1},a^{0})\sim\mu_{\mathcal{D}}(\cdot,\cdot,\cdot)}\Big{[}D_{\mathrm{Hellinger}}^{2}\big{(}\mathbb{P}_{r^{\star}}(\cdot|x,a^{1},a^{0})\|\mathbb{P}_{r}(\cdot|x,a^{1},a^{0})\big{)}\Big{]}$ (A.71)
$\qquad\qquad+\frac{3}{N}\cdot\log\left(\frac{\mathcal{N}_{\varepsilon}(\mathcal{R},\|\cdot\|_{\infty})}{\delta}\right).$ (A.72)
This completes the proof of Lemma A.1. ∎
Lemma A.2 (Sigmoid function).
@@
$\kappa\cdot|z_{1}-z_{2}|\leq\left|\sigma(z_{1})-\sigma(z_{2})\right|\leq|z_{1}-z_{2}|,$ (A.73)
@@
$\sigma(z_{1})-\sigma(z_{2})=\sigma^{\prime}\big{(}\xi(z_{1},z_{2})\big{)}\cdot(z_{1}-z_{2}).$ (A.74)
@@
$1\geq\sigma^{\prime}\big{(}\xi(z_{1},z_{2})\big{)}$ $=\sigma\big{(}\xi(z_{1},z_{2})\big{)}\cdot\Big{(}1-\sigma\big{(}\xi(z_{1},z_{2})\big{)}\Big{)}$ (A.75)
$=\frac{1}{1+\exp(\xi(z_{1},z_{2}))}\cdot\left(1-\frac{1}{1+\exp(\xi(z_{1},z_{2}))}\right)$ (A.76)
$\geq\frac{1}{1+\exp(R)}\cdot\left(1-\frac{1}{1+\exp(-R)}\right)$ (A.77)
$=\frac{1}{(1+\exp(R))^{2}}.$ (A.78)
This completes the proof of Lemma A.2. ∎
Appendix B Proofs for Equivalence between Maximin and Minimax Objectives
B.1 Proof of Theorem 5.6
@@ Appendix B Proofs for Equivalence between Maximin and Minimax Objectives
$\widehat{\pi}\in\mathop{\mathrm{argmax}}_{\pi\in\Pi}\,\min_{r\in\mathcal{R}}\phi(\pi,r).$ (B.1)
@@
$\mathrm{Dual}(\widehat{r},\widehat{\pi}):=\max_{\pi\in\Pi}\phi(\pi,\widehat{r})-\min_{r\in\mathcal{R}}\phi(\widehat{\pi},r)$ (B.2)
@@
$\mathrm{Dual}(\widehat{r},\widehat{\pi})$ $=\left(\max_{\pi\in\Pi}\phi(\pi,\widehat{r})-\min_{r\in\mathcal{R}}\,\max_{\pi\in\Pi}\phi(\pi,r)\right)$ (B.3)
$\qquad-\left(\max_{\pi\in\Pi}\,\min_{r\in\mathcal{R}}\phi(\pi,r)-\min_{r\in\mathcal{R}}\phi(\widehat{\pi},r)\right)$ (B.4)
$=0,$ (B.5)
where in the first equality we apply Lemma B.1 that the minimax objective and the maximin objective are equivalent, and the last equality applies the definition of $\widehat{r}$ and $\widehat{\pi}$ respectively. Note that we can rewrite the duality gap as following
$\mathrm{Dual}(\widehat{r},\widehat{\pi})=\left(\max_{\pi\in\Pi}\phi(\pi,\widehat{r})-\phi(\widehat{\pi},\widehat{r})\right)-\left(\phi(\widehat{\pi},\widehat{r})-\min_{r\in\mathcal{R}}\phi(\widehat{\pi},r)\right).$ (B.6)
Combining (B.5) and (B.6), we can conclude that
$\max_{\pi\in\Pi}\phi(\pi,\widehat{r})=\phi(\widehat{\pi},\widehat{r})\quad\Rightarrow\quad\widehat{\pi}\in\mathop{\mathrm{argmax}}_{\pi\in\Pi}\phi(\widehat{r},\pi).$ (B.7)
Now comparing what $\pi_{\widehat{r}}$ and $\widehat{\pi}$ satisfy in (5.7) and (B.7) respectively, invoking Lemma B.3 that the maximizer of $\phi(\cdot,r)$ given any $r\in\mathcal{R}$ is unique on the support of $d_{0}$, we can conclude that
$\pi_{\widehat{r}}(\cdot|x)=\widehat{\pi}(\cdot|x),\quad\forall x\in\mathrm{Supp}(d_{0}).$ (B.8)
Therefore, by (B.1) and (B.8), and the fact that $\phi(\pi,r)$ depends on $\pi$ only through its value on the support of $d_{0}$, we can conclude that
$\pi_{\widehat{r}}\in\mathop{\mathrm{argmax}}_{\pi\in\Pi}\,\min_{r\in\mathcal{R}}\phi(\pi,r).$ (B.9)
@@
B.2 Auxiliary Lemmas
Lemma B.1 (Equivalence of maximin and minimax objectives).

### v2

@@ B.2 Technical Lemmas
$\mathcal{L}_{\mathcal{D}}(r^{\star})-\mathcal{L}_{\mathcal{D}}(r)$ $\leq-2\cdot\mathbb{E}_{(x,a^{1},a^{0})\sim\mu_{\mathcal{D}}(\cdot,\cdot,\cdot)}\Big{[}D_{\mathrm{Hellinger}}^{2}\big{(}\mathbb{P}_{r^{\star}}(\cdot|x,a^{1},a^{0})\|\mathbb{P}_{r}(\cdot|x,a^{1},a^{0})\big{)}\Big{]}$ (B.47)
$\qquad+\frac{3}{N}\cdot\log\left(\frac{\mathcal{N}_{\varepsilon}(\mathcal{R},\|\cdot\|_{\infty})}{\delta}\right).$ (B.48)
Proof of Lemma B.1.
@@
$\mathcal{L}_{\mathcal{D}}(r^{\star})-\mathcal{L}_{\mathcal{D}}(r)$ $\leq-2\cdot\mathbb{E}_{(x,a^{1},a^{0})\sim\mu_{\mathcal{D}}(\cdot,\cdot,\cdot)}\Big{[}D_{\mathrm{Hellinger}}^{2}\big{(}\mathbb{P}_{r^{\star}}(\cdot|x,a^{1},a^{0})\|\mathbb{P}_{r}(\cdot|x,a^{1},a^{0})\big{)}\Big{]}$ (B.49)
$\qquad+\frac{2}{N}\cdot\log\left(\frac{\mathcal{N}_{\varepsilon}(\mathcal{R},\|\cdot\|_{\infty})}{\delta}\right).$ (B.50)
@@
$\mathcal{L}_{\mathcal{D}}(r^{\star})-\mathcal{L}_{\mathcal{D}}(r)$ $=\mathcal{L}_{\mathcal{D}}(r^{\star})-\mathcal{L}_{\mathcal{D}}(r^{\dagger})+\mathcal{L}_{\mathcal{D}}(r^{\dagger})-\mathcal{L}_{\mathcal{D}}(r)$ (B.51)
$\leq-2\cdot\mathbb{E}_{(x,a^{1},a^{0})\sim\mu_{\mathcal{D}}(\cdot,\cdot,\cdot)}\Big{[}D_{\mathrm{Hellinger}}^{2}\big{(}\mathbb{P}_{r^{\star}}(\cdot|x,a^{1},a^{0})\|\mathbb{P}_{r^{\dagger}}(\cdot|x,a^{1},a^{0})\big{)}\Big{]}$ (B.52)
$\qquad+\frac{2}{N}\cdot\log\left(\frac{\mathcal{N}_{\varepsilon}(\mathcal{R},\|\cdot\|_{\infty})}{\delta}\right)+\mathcal{L}_{\mathcal{D}}(r^{\dagger})-\mathcal{L}_{\mathcal{D}}(r)$ (B.53)
$\leq-2\cdot\mathbb{E}_{(x,a^{1},a^{0})\sim\mu_{\mathcal{D}}(\cdot,\cdot,\cdot)}\Big{[}D_{\mathrm{Hellinger}}^{2}\big{(}\mathbb{P}_{r^{\star}}(\cdot|x,a^{1},a^{0})\|\mathbb{P}_{r}(\cdot|x,a^{1},a^{0})\big{)}\Big{]}$ (B.54)
$\qquad+\frac{2}{N}\cdot\log\left(\frac{\mathcal{N}_{\varepsilon}(\mathcal{R},\|\cdot\|_{\infty})}{\delta}\right)+\mathcal{L}_{\mathcal{D}}(r^{\dagger})-\mathcal{L}_{\mathcal{D}}(r)$ (B.55)
$\qquad\qquad+4\cdot\mathbb{E}_{(x,a^{1},a^{0})\sim\mu_{\mathcal{D}}(\cdot,\cdot,\cdot)}\Big{[}D_{\mathrm{Hellinger}}^{2}\big{(}\mathbb{P}_{r^{\dagger}}(\cdot|x,a^{1},a^{0})\|\mathbb{P}_{r}(\cdot|x,a^{1},a^{0})\big{)}\Big{]},$ (B.56)
where in the fir inequality we use (B.50) for $r^{\dagger}$ and in the second inequality we utilize the triangular inequality for Hellinger distance. Therefore, it remains to upper bound the approximation error induced by $r^{\dagger}$. On the one hand, by the definition of $\mathcal{L}_{\mathcal{D}}$ in (3.2), we have that
$\mathcal{L}_{\mathcal{D}}(r^{\dagger})-\mathcal{L}_{\mathcal{D}}(r)$ $=\frac{1}{N}\sum_{i=1}^{N}y_{i}\cdot\log\left(\frac{\sigma\big{(}r(x_{i},a^{1}_{i})-r(x_{i},a^{0}_{i})\big{)}}{\sigma\big{(}r^{\dagger}(x_{i},a^{1}_{i})-r^{\dagger}(x_{i},a^{0}_{i})\big{)}}\right)$ (B.57)
$\qquad+\frac{1}{N}\sum_{i=1}^{N}(1-y_{i})\cdot\log\left(\frac{\sigma\big{(}r(x_{i},a^{0}_{i})-r(x_{i},a^{1}_{i})\big{)}}{\sigma\big{(}r^{\dagger}(x_{i},a^{0}_{i})-r^{\dagger}(x_{i},a^{1}_{i})\big{)}}\right).$ (B.58)
@@
$\mathcal{L}_{\mathcal{D}}(r^{\dagger})-\mathcal{L}_{\mathcal{D}}(r)$ $\leq\frac{1}{N}\sum_{i=1}^{N}y_{i}\cdot\frac{\sigma\big{(}r(x_{i},a^{1}_{i})-r(x_{i},a^{0}_{i})\big{)}-\sigma\big{(}r^{\dagger}(x_{i},a^{1}_{i})-r^{\dagger}(x_{i},a^{0}_{i})\big{)}}{\sigma\big{(}r^{\dagger}(x_{i},a^{1}_{i})-r^{\dagger}(x_{i},a^{0}_{i})\big{)}}$ (B.59)
$\qquad\qquad+\frac{1}{N}\sum_{i=1}^{N}(1-y_{i})\cdot\frac{\sigma\big{(}r(x_{i},a^{0}_{i})-r(x_{i},a^{1}_{i})\big{)}-\sigma\big{(}r^{\dagger}(x_{i},a^{0}_{i})-r^{\dagger}(x_{i},a^{1}_{i})\big{)}}{\sigma\big{(}r^{\dagger}(x_{i},a^{0}_{i})-r^{\dagger}(x_{i},a^{1}_{i})\big{)}}.$ (B.60)
Now since $\|r^{\dagger}-r\|_{\infty}\leq\varepsilon$ and $r^{\dagger}\in[0,R]$, invoking Lemma B.2, we can derive that
$\mathcal{L}_{\mathcal{D}}(r^{\dagger})-\mathcal{L}_{\mathcal{D}}(r)$ $\leq\frac{1}{N}\sum_{i=1}^{N}\frac{\big{|}\big{(}r(x_{i},a^{1}_{i})-r(x_{i},a^{0}_{i})\big{)}-\big{(}r^{\dagger}(x_{i},a^{1}_{i})-r^{\dagger}(x_{i},a^{0}_{i})\big{)}\big{|}}{(1+e^{R})^{-1}}$ (B.61)
$\qquad+\frac{1}{N}\sum_{i=1}^{N}\frac{\big{|}\big{(}r(x_{i},a^{0}_{i})-r(x_{i},a^{1}_{i})\big{)}-\big{(}r^{\dagger}(x_{i},a^{0}_{i})-r^{\dagger}(x_{i},a^{1}_{i})\big{)}\big{|}}{(1+e^{R})^{-1}}$ (B.62)
$\leq 4\cdot\|r^{\dagger}-r\|_{\infty}\cdot(1+e^{R})\leq 4\varepsilon\cdot(1+e^{R}).$ (B.63)
@@
$D_{\mathrm{Hellinger}}^{2}\big{(}\mathbb{P}_{r^{\dagger}}(\cdot|x,a^{1},a^{0})\|\mathbb{P}_{r}(\cdot|x,a^{1},a^{0})\big{)}$ (B.64)
$\qquad\leq D_{\mathrm{TV}}\big{(}\mathbb{P}_{r^{\dagger}}(\cdot|x,a^{1},a^{0})\|\mathbb{P}_{r}(\cdot|x,a^{1},a^{0})\big{)}$ (B.65)
$\qquad=\Big{|}\sigma\big{(}r^{\dagger}(x,a^{1})-r^{\dagger}(x,a^{0})\big{)}-\sigma\big{(}r(x,a^{1})-r(x,a^{0})\big{)}\Big{|}$ (B.66)
$\qquad\leq\Big{|}\big{(}r^{\dagger}(x,a^{1})-r^{\dagger}(x,a^{0})\big{)}-\big{(}r(x,a^{1})-r(x,a^{0})\big{)}\Big{|}$ (B.67)
$\qquad\leq 2\cdot\|r^{\dagger}-r\|_{\infty}\leq 2\varepsilon,$ (B.68)
where the first inequality uses the fact that $D_{\mathrm{Hellinger}}^{2}\leq D_{\mathrm{TV}}$, the equality uses the same argument as (B.35), and the second inequality applies Lemma B.2. Finally, combining (B.56), (B.63), and (B.68), we conclude that
$\mathcal{L}_{\mathcal{D}}(r^{\star})-\mathcal{L}_{\mathcal{D}}(r)$ $\leq-2\cdot\mathbb{E}_{(x,a^{1},a^{0})\sim\mu_{\mathcal{D}}(\cdot,\cdot,\cdot)}\Big{[}D_{\mathrm{Hellinger}}^{2}\big{(}\mathbb{P}_{r^{\star}}(\cdot|x,a^{1},a^{0})\|\mathbb{P}_{r}(\cdot|x,a^{1},a^{0})\big{)}\Big{]}$ (B.69)
$\qquad+\frac{2}{N}\cdot\log\left(\frac{\mathcal{N}_{\varepsilon}(\mathcal{R},\|\cdot\|_{\infty})}{\delta}\right)+6\varepsilon\cdot(1+e^{R}).$ (B.70)
@@
$\mathcal{L}_{\mathcal{D}}(r^{\star})-\mathcal{L}_{\mathcal{D}}(r)$ $\leq-2\cdot\mathbb{E}_{(x,a^{1},a^{0})\sim\mu_{\mathcal{D}}(\cdot,\cdot,\cdot)}\Big{[}D_{\mathrm{Hellinger}}^{2}\big{(}\mathbb{P}_{r^{\star}}(\cdot|x,a^{1},a^{0})\|\mathbb{P}_{r}(\cdot|x,a^{1},a^{0})\big{)}\Big{]}$ (B.71)
$\qquad\qquad+\frac{3}{N}\cdot\log\left(\frac{\mathcal{N}_{\varepsilon}(\mathcal{R},\|\cdot\|_{\infty})}{\delta}\right).$ (B.72)
This completes the proof of Lemma B.1. ∎
Lemma B.2 (Sigmoid function).
@@
$\kappa\cdot|z_{1}-z_{2}|\leq\left|\sigma(z_{1})-\sigma(z_{2})\right|\leq|z_{1}-z_{2}|,$ (B.73)
@@
$\sigma(z_{1})-\sigma(z_{2})=\sigma^{\prime}\big{(}\xi(z_{1},z_{2})\big{)}\cdot(z_{1}-z_{2}).$ (B.74)
@@
$1\geq\sigma^{\prime}\big{(}\xi(z_{1},z_{2})\big{)}$ $=\sigma\big{(}\xi(z_{1},z_{2})\big{)}\cdot\Big{(}1-\sigma\big{(}\xi(z_{1},z_{2})\big{)}\Big{)}$ (B.75)
$=\frac{1}{1+\exp(\xi(z_{1},z_{2}))}\cdot\left(1-\frac{1}{1+\exp(\xi(z_{1},z_{2}))}\right)$ (B.76)
$\geq\frac{1}{1+\exp(R)}\cdot\left(1-\frac{1}{1+\exp(-R)}\right)$ (B.77)
$=\frac{1}{(1+\exp(R))^{2}}.$ (B.78)
This completes the proof of Lemma B.2. ∎
Appendix C Proofs for Equivalence between Maximin and Minimax Objectives
C.1 Proof of Theorem 5.6
@@ Appendix C Proofs for Equivalence between Maximin and Minimax Objectives
$\widehat{\pi}\in\mathop{\mathrm{argmax}}_{\pi\in\Pi}\,\min_{r\in\mathcal{R}}\phi(\pi,r).$ (C.1)
@@
$\mathrm{Dual}(\widehat{r},\widehat{\pi}):=\max_{\pi\in\Pi}\phi(\pi,\widehat{r})-\min_{r\in\mathcal{R}}\phi(\widehat{\pi},r)$ (C.2)
@@
$\mathrm{Dual}(\widehat{r},\widehat{\pi})$ $=\left(\max_{\pi\in\Pi}\phi(\pi,\widehat{r})-\min_{r\in\mathcal{R}}\,\max_{\pi\in\Pi}\phi(\pi,r)\right)$ (C.3)
$\qquad+\left(\max_{\pi\in\Pi}\,\min_{r\in\mathcal{R}}\phi(\pi,r)-\min_{r\in\mathcal{R}}\phi(\widehat{\pi},r)\right)$ (C.4)
$=0,$ (C.5)
where in the first equality we apply Lemma C.1 that the minimax objective and the maximin objective are equivalent, and the last equality applies the definition of $\widehat{r}$ and $\widehat{\pi}$ respectively. Note that we can rewrite the duality gap as following
$\mathrm{Dual}(\widehat{r},\widehat{\pi})=\left(\max_{\pi\in\Pi}\phi(\pi,\widehat{r})-\phi(\widehat{\pi},\widehat{r})\right)+\left(\phi(\widehat{\pi},\widehat{r})-\min_{r\in\mathcal{R}}\phi(\widehat{\pi},r)\right).$ (C.6)
Combining (C.5) and (C.6), we can conclude that
$\max_{\pi\in\Pi}\phi(\pi,\widehat{r})=\phi(\widehat{\pi},\widehat{r})\quad\Rightarrow\quad\widehat{\pi}\in\mathop{\mathrm{argmax}}_{\pi\in\Pi}\phi(\widehat{r},\pi).$ (C.7)
Now comparing what $\pi_{\widehat{r}}$ and $\widehat{\pi}$ satisfy in (5.7) and (C.7) respectively, invoking Lemma C.3 that the maximizer of $\phi(\cdot,r)$ given any $r\in\mathcal{R}$ is unique on the support of $d_{0}$, we can conclude that
$\pi_{\widehat{r}}(\cdot|x)=\widehat{\pi}(\cdot|x),\quad\forall x\in\mathrm{Supp}(d_{0}).$ (C.8)
Therefore, by (C.1) and (C.8), and the fact that $\phi(\pi,r)$ depends on $\pi$ only through its value on the support of $d_{0}$, we can conclude that
$\pi_{\widehat{r}}\in\mathop{\mathrm{argmax}}_{\pi\in\Pi}\,\min_{r\in\mathcal{R}}\phi(\pi,r).$ (C.9)
@@
C.2 Auxiliary Lemmas
Lemma C.1 (Equivalence of maximin and minimax objectives).